# Straggler-Resilient Decentralized Learning via Adaptive Asynchronous Updates

### v1

@@ 5 Realization of DSGD-AAU
Table 1: Test accuracy for different models on non-i.i.d. CIFAR-10 with 128 workers.
6 Numerical Results
We conduct extensive experiments to validate our model and theoretical results. We implement DSGD-AAU in PyTorch [paszke2017automatic] on Python 3 with three NVIDIA RTX A6000 GPUs using the Network File System (NFS) and MPI backends, where files are shared among workers through NFS and communications between workers are via MPI. We relegate some experimental details, parameter settings and a comprehensive set of results to the appendices.
Baseline. We compare DSGD-AAU with three state-of-the-art methods: AD-PSGD [lian2018asynchronous], Prague [luo2020prague] and AGP [assran2020asynchronous], where workers are randomly selected for parameter averaging in AD-PSGD and Prague. Hence both suffer from stragglers in the optimization problem since they ignore slower workers.
Dataset and Model. We consider four representative models: a fully-connected neural network with 2 hidden layers (2-NN), AlexNet [krizhevsky2012imagenet], VGG-13 [simonyan2015very] and ResNet-18 [he2016deep] for classification tasks using CIFAR-10 [krizhevsky2009learning], MNIST [lecun1998gradient], and Tiny-ImageNet [russakovsky2015imagenet] datasets, in particular focusing on non-i.i.d. versions. For example, we use the idea in [mcmahan2017communication] to obtain non-i.i.d. CIFAR-10. We also investigate the task of next-character prediction on the dataset of The Complete Works of William Shakespeare [mcmahan2017communication] (Shakespeare) with an LSTM language model [kim2016character]. We consider a network with $32$, $64$, $128$, $256$ workers and randomly generate a connected graph for evaluation. We randomly select workers as stragglers in each iteration, e.g., the straggler sleeps for some time in the iteration (e.g., the sleep time could be 6x of the average one local computation time), as suggested by [cipar2013solving, lian2018asynchronous, luo2020prague, li2020taming]. The loss function is the cross-entropy one.
@@ 6 Numerical Results
The learning rate is perhaps the most critical hyperparameter in distributed ML optimization problems. A proper value of the learning rate is important; however, it is in general hard to find the optimal value. With the consideration of adaptive asynchronous updates, it is reasonable to adaptively set the learning rate to speed up the convergence process in experiments, as in [chen2016revisiting, lian2018asynchronous, xu2020dynamic, assran2020asynchronous]. Hence, we choose $\eta(k)=\eta_{0}\cdot\delta^{k}$, where $\eta_{0}=0.1$ and $\delta=0.95.$ Finally, batch size is another important hyperparameter, which is limited by the memory and computational resources available at each worker, or determined by generalization performance of final model [hoffer2017train]. We test the impact of batch size using these datasets and find that $128$ is a proper value. An ablation study on these hyperparameters (e.g. straggler rate) is given in the appendices.
Figure 4: Training loss w.r.t. time for different models on non-i.i.d. CIFAR-10 with 128 workers.
@@
Convergence. Figure 3 shows the training loss for all algorithms w.r.t. iterations, which are evaluated for different models on non-i.i.d. CIFAR-10 dataset with $128$ workers. The corresponding test accuracy is presented in Table 1. We observe that DSGD-AAU significantly outperforms AD-PSGD and AGP, and performs better than Prague when the total number of iterations is large enough. Its advantage is especially pronounced with a limited number of iterations or a limited training time (see Figure 4), e.g., within $6,000$ iterations, DSGD-AAU achieves much smaller loss (resp. higher test accuracy) than Prague for ResNet-18. Finally, note that the test accuracy for non-i.i.d. CIFAR-10 in Table 1 is relatively lower compared to that for i.i.d. dataset, a phenomenon widely observed in the literature [yang2021achieving, zhao2018federated, yeganeh2020inverse, wang2020optimizing, hsieh2020non].
Similar observations can be made with $32$, $64$, $256$ workers across other datasets and models (see the appendices). In particular, we summarize the test accuracy for all algorithms for ResNet-18 on non-i.i.d. CIFAR-10 trained for $50$ seconds with different number of workers in Table 2. DSGD-AAU also achieves a higher test accuracy compared to AD-PSGD, Prague and AGP.
@@
Table 2: Test accuracy for ResNet-18 on non-i.i.d. CIFAR-10 trained for $50$ seconds with different number of workers.
Speedup and Communication. The convergence in wall-clock time of ResNet-18 on non-i.i.d. CIFAR-10 dataset with $128$ workers is presented in Figure 4. Again, we observe that DSGD-AAU converges faster, and importantly, DSGD-AAU achieves a higher accuracy for a given training time. The speedup of different algorithms for ResNet-18 on non-i.i.d. CIFAR-10 with respect to different number of workers is presented in Figure 5(a). We compute the speedup of different algorithms with respect to the DSGD with full worker updates when $65\%$ accuracy is achieved for ResNet-18 model. DSGD-AAU consistently outperforms baseline methods and achieves the best speedup at no cost of additional communication (i.e., network transmission for training) in the system as shown in Figure 5(b). Similar observations can be made for other models with different datasets (see the appendices), i.e., DSGD-AAU is robust across different models and datasets to achieve the fastest convergence time and the best speedup while maintaining comparable or even better communication efficiency.
@@
Figure 5: Speedup and communication for ResNet-18 on non-i.i.d. CIFAR-10 with different number of workers.
7 Related Work
@@
Our algorithm belongs to the class of consensus/gossip algorithms, or fully decentralized/peer-to-peer learning algorithms [kairouz2019advances, yang2019federated, imteaj2021survey, imteaj2022federated]. There is a rich literature of consensus algorithms based on the theory of [tsitsiklis1986distributed] in various fields including [bertsekas1989parallel, kempe2003gossip, xiao2004fast, boyd2006randomized, nedic2009distributed, johansson2010randomized, ram2010distributed, boyd2011distributed, duchi2011dual, colin2016gossip, scaman2017optimal, scaman2018optimal, tang2018communication, tang2018d, li2019consensus, neglia2020decentralized]. However, most results are for full worker participation with synchronous updates in each iteration, which can be communication expensive. Efficient control of adaptive asynchronous updates in consensus-based distributed optimization has received little attention, which arise in problems of practical interests. The few existing works address this issue through a stale-synchronous model [chen2016revisiting, ho2013more] with no convergence guarantee, and the number of fastest neighbor workers is often configured manually through preliminary experiments before the start of the actual training process.
@@
The asynchronous SGD [agarwal2011distributed] breaks the synchronization in synchronous SGD by allowing workers to use stale weights to compute gradients, which significantly reduces the idel time for synchronization. [lian2018asynchronous] proposed the first asynchronous decentralized parallel SGD algorithm AD-PSGD which achieves linear speedup with the number of workers. The limitation of AD-PSGD is that the network topology needs be bipartite to avoid deadlock issue. [luo2020prague] proposed Prague based on a slight modification of AD-PSGD, which resolves the deadlock issue and further reduces the convergence times. [assran2020asynchronous] proposed asynchronous gradient push (AGP) to deal with directed graph with non-doubly stochastic consensus matrix under the assumption that the loss function is convex. A similar gradient-push technique was used for decentralized online learning in [jiang2021asynchronous]. For non-convex stochastic optimization on directed graphs, [kungurtsev2021decentralized] proposed an asynchronous algorithm which combines stochastic gradients with tracking in an asynchronous push-sum framework and obtains the standard sublinear convergence rate. [nadiradze2021asynchronous] proposed an asynchronous decentralized SGD with quantized local updates to reduce the communication overhead. [wang2022asynchronous] proposed a predicting Clipping asynchronous SGD where the predicting step leverages the gradient prediction using Taylor expansion to reduce the staleness of the outdated weights while the clipping step selectively drops the outdated weights to alleviate their negative effects. More related work can be found in [assran2020advances]. A more detailed review of fully decentralized learning algorithms as well as the All-Reduce [li2020taming, li2020breaking], or the parameter-server paradigm [karakus2017straggler, charalambides2020weighted, glasgow2021approximate], such as the popular federated learning [mcmahan2017communication], is provided in the appendices.
8 Conclusion
@@
Appendix A Additional Related Work
We provide additional discussions on the related work. In particular, we compare our DSGD-AAU with two state-of-the-arts AD-PSGD [lian2018asynchronous] and Prague [luo2020prague]. Similar to traditional training such as PS and All-Reduce, in one iteration, AD-PSGD computes gradients first, and then performs synchronization; the difference is that it only synchronizes with a randomly selected neighbor, instead of all other workers. However, a conflict will occur in AD-PSGD when two workers simultaneously select one worker for synchronization. In order to keep the atomic property of weight updating, the two operations need to be serialized, which slows down the whole process. To resolve the conflict issue and fasten the synchronization, Prague proposed a partial All-Reduce scheme which randomly generates a group for each node, and the node only communicates with other nodes in the same group. However, the performance of Prague heavily depends on a so-called “Group Generator”, i.e., when one worker completes its local update, it would inquire the Group Generator to get its group number, and then it has to wait all members in the corresponding group to complete their local updates before implementing the communication process (i.e., sharing the local updates with the group members). In particular, three methods were proposed in [luo2020prague] including an offline static method, a randomized method and a “smart” method using a heuristic threshold-based rule. However, the method of how to adjust the threshold was not discussed.
In addition, a similar idea of adaptive asynchronous update has been widely used in DML systems, there exist significant challenges. Previous works mostly focused on the PS or the All-Reduce models, and most of existing algorithms are network topology-constrained and are hard to be generalized and implemented in arbitrary network topologies as our proposed DSGD-AAU does. Moreover, most algorithms including in Li et al. [li2020taming, li2020breaking] only select the fastest few workers for the parameter averaging process, where a lot of “normal” workers are ignored to drag down the whole update process. Furthermore, another line of works using coding techniques such as [karakus2017straggler, charalambides2020weighted, glasgow2021approximate] aim to recover the stragglers’ gradient information by leveraging gradient coding techniques from information theory perspective. However, same data needs to be sent to multiple nodes to enable gradient encoding in these works, while the dataset for each node is disjoint in our model. Hence each node in our model has non-i.i.d. distinct dataset, while cannot be applied in [karakus2017straggler, charalambides2020weighted, glasgow2021approximate].
A more relevant line of research on distributed learning is the All-Reduce [li2020taming, li2020breaking], or the PS [karakus2017straggler, charalambides2020weighted, glasgow2021approximate]. In particular, the popular federated learning [mcmahan2017communication] belongs to the PS paradigm which has received increased attentions in recent years. We refer interested readers to [kairouz2019advances] and references therein for a complete review. Most existing work is based on the centralized framework through synchronous communication scheme which often leads to communication congestion in the server. Alternatively, the PS can operate asynchronously, updating parameter vector immediately after it receives updates from a single worker. Although such an asynchronous update manner can increase system throughput, some workers may still operate on stale versions of the parameter vector and in some cases, even preventing convergence to the optimal model [dai2018toward]. More recently, several PS-based algorithms considering stale-synchronous model via adaptive worker participation have gained increased attention. See [chen2016revisiting, teng2018bayesian, luo2019hop, ruan2020towards, xu2020dynamic, yang2021achieving] for details, among which [xu2020dynamic] is perhaps the only work proposing to adaptively determine the number of fastest neighbor workers. However, most of them focus on showing empirical results without convergence guarantees and without a rigorous analysis of how selection skew affects convergence speed. Moreover, none of them exploit the communication among workers, since they assume either a central server exists to coordinate workers (e.g., PS model) or all workers have an identical role (e.g., All-Reduce). Generalizing the PS schemes to our fully decentralized framework is a non-trivial and open problem because of the large scale distributed and heterogeneous nature of the training data.
Appendix B The Proposed Path Searching Procedure
@@
Algorithm 3 Pathsearch -logical view of arbitrary worker
Appendix C Proofs of Main Result
In this section, we provide proofs of the theoretical results presented in the paper, i.e., the convergence for DSGD-AAU after a large number of iterations (Theorem 1).
C.1 Proof of Theorem 1
@@
and let ${\Phi}_{k,s}(i,j)$ be the $i$-th row $j$-th column element of $\boldsymbol{\Phi}_{k:s}$. Therefore, with fixed $\eta$, ${\mathbf{w}}_{i}(k)$ can be expressed in terms of $\boldsymbol{\Phi}$ as
${\mathbf{w}}_{i}(k+1)$ $=\sum\limits_{j=1}^{N}\Phi_{k,1}(i,j){\mathbf{w}}_{j}(0)-\eta\sum\limits_{r=1}^{k}\sum\limits_{j=1}^{N}\Phi_{k,r}(i,j)g_{i}({\mathbf{w}}_{j}(r-1)).$
@@
$\bar{{\mathbf{w}}}(k)=\frac{1}{N}\sum\limits_{j=1}^{N}{\mathbf{w}}_{j}(0)-\eta\sum\limits_{r=1}^{k}\sum\limits_{j=1}^{N}\frac{1}{N}g({\mathbf{w}}_{j}(r-1)),$ (8)
@@
$\bar{{\mathbf{w}}}(k+1)=\bar{{\mathbf{w}}}(k)-\frac{\eta}{N}\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k)).$ (9)
Due to the gradient Lipschitz-continuous in Assumption 4.3, we have the following inequality with respect to $\bar{{\mathbf{w}}}(k)$ [bottou2018optimization]
@@
$\qquad\qquad+\frac{L}{2}\mathbb{E}\left[\|\bar{{\mathbf{w}}}(k+1)\!-\!\bar{{\mathbf{w}}}(k)\|^{2}\right]\!+\!\mathbb{E}[F(\bar{{\mathbf{w}}}(k))]$
@@
$\qquad\qquad+\underset{C_{2}}{\underbrace{\frac{L}{2}\mathbb{E}\left[\|\bar{{\mathbf{w}}}(k+1)-\bar{{\mathbf{w}}}(k)\|^{2}\right]}}\!+\!\mathbb{E}[F(\bar{{\mathbf{w}}}(k))]\!-\!\eta\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}.$ (10)
As a result, we need to bound $C_{1}$ and $C_{2}$ in (C.1). The term $C_{1}$ can be bounded as follows
@@
$\qquad\qquad-\frac{\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))\right\|^{2}+\frac{\eta}{2}\left\|\nabla F(\bar{{\mathbf{w}}}(k))\right\|^{2}$
@@
$\qquad\qquad-\frac{\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))\right\|^{2}+\frac{\eta}{2}\left\|\nabla F(\bar{{\mathbf{w}}}(k))\right\|^{2},$ (11)
where $(a_{1})$ follows from (9) and the fact that $\frac{1}{N}\sum\limits_{j=1}^{N}\nabla F_{j}(\bar{{\mathbf{w}}}(k))=\nabla F(\bar{{\mathbf{w}}}(k))$; $(a_{2})$ comes from a standard mathematical manipulation; the equality $(a_{3})$ is based on the equation $\langle{\mathbf{a}},{\mathbf{b}}\rangle=\frac{1}{2}\left[\|{\mathbf{a}}\|^{2}+\|{\mathbf{b}}\|^{2}-\|{\mathbf{a}}-{\mathbf{b}}\|^{2}\right]$ and $\mathbb{E}[g_{j}({\mathbf{w}}_{j}(k))]=\nabla F_{j}({\mathbf{w}}_{j}(k))$ from Assumption 4.4; and the last inequality $(a_{4})$ is due to the well triangle-inequality $\left\|\sum\limits_{i=1}^{N}{\mathbf{a}}_{i}\right\|^{2}\leq N\sum\limits_{i}^{N}\left\|{\mathbf{a}}_{i}\right\|^{2}.$
@@
$D_{1}{\leq}$ $\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\Big{[}\|{\mathbf{w}}_{j}(k)-\bar{{\mathbf{w}}}(k)\|^{2}\Big{]}$
where the inequality directly comes from Assumption 4.3.
Substituting $D_{1}$ into $C_{1}$ yields
$C_{1}$ $\leq\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\left\|{\mathbf{w}}_{j}(k)-\bar{{\mathbf{w}}}(k)\right\|^{2}-\frac{\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))\right\|^{2}+\frac{\eta}{2}\left\|\nabla F(\bar{{\mathbf{w}}}(k))\right\|^{2}.$ (12)
@@
=$\overset{(c_{1})}{=}\frac{\eta^{2}L}{2N^{2}}\left(\mathbb{E}\left[\left\|\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))-\nabla F_{j}({\mathbf{w}}_{j}(k))\right\|^{2}\right]+\left\|\mathbb{E}\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))\right\|^{2}\right)$
$\overset{(c_{2})}{\leq}\frac{\eta^{2}L}{2N}\sigma^{2}_{L}+\frac{\eta^{2}L}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))\right\|^{2},$ (13)
@@
Substituting $C_{1}$ and $C_{2}$ into the inequality (C.1), we have

### v2

@@ 6. Numerical Results
Table 1. Test accuracy for different models on non-i.i.d. CIFAR-10 with 128 workers.
@@
The loss function is the cross-entropy one. The learning rate is a critical hyperparameter in distributed ML optimization problems. A proper value is important; however, it is in general hard to find the optimal value. With the consideration of adaptive asynchronous updates, it is reasonable to adaptively set the learning rate to speed up the convergence process in experiments (Chen et al., 2016; Lian et al., 2018; Xu et al., 2020; Assran and Rabbat, 2020). Hence, we choose $\eta(k)=\eta_{0}\cdot\delta^{k}$, where $\eta_{0}=0.1$ and $\delta=0.95.$ Finally, batch size is often limited by the memory and computational resources available at each worker, or determined by generalization performance of final model (Hoffer et al., 2017). We test the impact of batch size using these datasets and find that $128$ is a proper value. For ease of readability, in this section, we only present the numerical results on non-i.i.d. partitoned CIFAR-10, and relegate the corresponding results on other datasets to Appendix C.
@@
Convergence. Figure 3 shows the training loss for all algorithms w.r.t. iterations, which are evaluated for different models on non-i.i.d. CIFAR-10 dataset with $128$ workers. The corresponding test accuracy is presented in Table 1. We observe that DSGD-AAU significantly outperforms AD-PSGD and AGP, and performs better than Prague when the total number of iterations is large enough. Its advantage is especially pronounced with a limited number of iterations or a limited training time (see Figure 4), e.g., within $6,000$ iterations, DSGD-AAU achieves much smaller loss (resp. higher test accuracy) than Prague for ResNet-18. Finally, note that the test accuracy for non-i.i.d. CIFAR-10 in Table 1 is relatively lower compared to that for i.i.d. dataset, a phenomenon widely observed in the literature (Yang et al., 2021; Zhao et al., 2018; Yeganeh et al., 2020; Wang et al., 2020a; Hsieh et al., 2020). Similar observations can be made with $32$, $64$, $256$ workers across other datasets and models. In particular, we summarize the test accuracy for all algorithms for ResNet-18 on non-i.i.d. CIFAR-10 trained for $50$ seconds with different number of workers in Table 2. DSGD-AAU also achieves a higher test accuracy compared to AD-PSGD, Prague and AGP.
@@
Table 2. Test accuracy for ResNet-18 on non-i.i.d. CIFAR-10 trained for $50$ seconds with different number of workers.
[…]
@@
Figure 5. Speedup and communication for ResNet-18 on non-i.i.d. CIFAR-10 with different number of workers.
7. Related Work
@@
Our algorithm belongs to the class of consensus/gossip algorithms, or fully decentralized/peer-to-peer learning algorithms (Kairouz et al., 2021; Imteaj et al., 2021, 2022). There is a rich literature of consensus algorithms based on the theory of (Tsitsiklis et al., 1986) in various fields including (Boyd et al., 2006; Nedic and Ozdaglar, 2009; Johansson et al., 2010; Ram et al., 2010; Boyd et al., 2011; Colin et al., 2016; Scaman et al., 2017, 2018; Tang et al., 2018a, b; Li et al., 2019; Neglia et al., 2020; Xiong et al., 2024). However, most results are for full worker participation with synchronous updates in each iteration, which can be communication expensive. Efficient control of adaptive asynchronous updates in consensus-based distributed optimization has received little attention, which arise in problems of practical interests. The few existing works address this issue through a stale-synchronous model (Chen et al., 2016; Ho et al., 2013) with no convergence guarantee, and the number of fastest neighbor workers is configured manually through preliminary experiments before the start of training process.
@@
Asynchronous SGD (Agarwal and Duchi, 2011) breaks the synchronization in synchronous SGD by allowing workers to use stale weights to compute gradients, which significantly reduces the idel time for synchronization. (Lian et al., 2018) proposed the first asynchronous decentralized parallel SGD algorithm AD-PSGD which achieves linear speedup with the number of workers. The limitation of AD-PSGD is that the network topology needs be bipartite to avoid deadlock issue. (Luo et al., 2020) proposed Prague based on a slight modification of AD-PSGD, which resolves the deadlock issue and further reduces the convergence times. (Assran and Rabbat, 2020) proposed asynchronous gradient push to deal with directed graph with non-doubly stochastic consensus matrix under the assumption that the loss function is convex. A similar gradient-push technique was used for decentralized online learning in (Jiang et al., 2021). For non-convex stochastic optimization on directed graphs, (Kungurtsev et al., 2021) proposed an asynchronous algorithm combining stochastic gradients with tracking in an asynchronous push-sum framework and obtains a sublinear convergence rate. (Nadiradze et al., 2021) proposed an asynchronous decentralized SGD with quantized local updates to reduce communication overhead. (Wang et al., 2022) proposed a predicting Clipping asynchronous SGD where the predicting step leverages the gradient prediction using Taylor expansion to reduce the staleness of the outdated weights while the clipping step selectively drops the outdated weights to alleviate their negative effects. More related work can be found in (Assran et al., 2020).
Further discussions on differences with most related works. In particular, we compare our DSGD-AAU with two state-of-the-arts AD-PSGD (Lian et al., 2018) and Prague (Luo et al., 2020). Similar to traditional training such as PS and All-Reduce, in one iteration, AD-PSGD computes gradients first, and then performs synchronization; the difference is that it only synchronizes with a randomly selected neighbor, instead of all other workers. However, a conflict will occur in AD-PSGD when two workers simultaneously select one worker for synchronization. In order to keep the atomic property of weight updating, the two operations need to be serialized, which slows down the whole process. To resolve the conflict issue and fasten the synchronization, Prague proposed a partial All-Reduce scheme which randomly generates a group for each node, and the node only communicates with other nodes in the same group. However, the performance of Prague heavily depends on a so-called “Group Generator”, i.e., when one worker completes its local update, it would inquire the Group Generator to get its group number, and then it has to wait all members in the corresponding group to complete their local updates before implementing the communication process (sharing local updates with the group members). In particular, three methods were proposed in (Luo et al., 2020) including an offline static method, a randomized method and a “smart” method using a heuristic threshold-based rule. However, the method of how to adjust the threshold was not discussed.
All-Reduce and the parameter-server paradigm (Federated learning). In addition, a similar idea of adaptive asynchronous update has been widely used in DML systems, there exist significant challenges. Previous works mostly focused on the PS or the All-Reduce models, and most of existing algorithms are network topology-constrained and are hard to be generalized and implemented in arbitrary network topologies as our proposed DSGD-AAU does. Moreover, most algorithms including in Li et al. (Li et al., 2020a, b) only select the fastest few workers for the parameter averaging process, where a lot of “normal” workers are ignored to drag down the whole update process. Furthermore, another line of works using coding techniques such as (Karakus et al., 2017; Charalambides et al., 2020; Glasgow and Wootters, 2021) aim to recover the stragglers’ gradient information by leveraging gradient coding techniques from information theory perspective. However, same data needs to be sent to multiple nodes to enable gradient encoding in these works, while the dataset for each node is disjoint in our model. Hence each node in our model has non-i.i.d. distinct dataset, while cannot be applied in (Karakus et al., 2017; Charalambides et al., 2020; Glasgow and Wootters, 2021).
A more relevant line of research on distributed learning is the All-Reduce (Li et al., 2020a, b), or the PS (Karakus et al., 2017; Charalambides et al., 2020; Glasgow and Wootters, 2021; Yan et al., 2023). In particular, the popular federated learning (McMahan et al., 2017) belongs to the PS paradigm. We refer interested readers to (Kairouz et al., 2021) for a complete review. Most existing work is based on the centralized framework through synchronous communication scheme which often leads to communication congestion in the server. Alternatively, the PS can operate asynchronously, updating parameter vector immediately after it receives updates from a single worker. Although such an asynchronous update manner can increase system throughput, some workers may still operate on stale versions of the parameter vector and in some cases, even preventing convergence to the optimal model (Dai et al., 2018). More recently, several PS-based algorithms considering stale-synchronous model via adaptive worker participation have gained increased attention. See (Chen et al., 2016; Teng and Wood, 2018; Luo et al., 2019; Xu et al., 2020; Yang et al., 2021) for details, among which (Xu et al., 2020) is perhaps the only work proposing to adaptively determine the number of fastest neighbor workers. However, most of them focus on showing empirical results without convergence guarantees and without a rigorous analysis of how selection skew affects convergence speed. Moreover, none of them exploit the communication among workers, since they assume either a central server exists to coordinate workers (e.g., PS model) or all workers have an identical role (e.g., All-Reduce). Generalizing the PS schemes to our fully decentralized framework is a non-trivial and open problem because of the large scale distributed and heterogeneous nature of the training data.
Communication stragglers. Although we focus on stragglers caused by heterogeneous computation capabilities among workers as in previous work in (Ananthanarayanan et al., 2013; Ho et al., 2013; Lian et al., 2017, 2018; Luo et al., 2020),there is another line of works focusing on another key type of stragglers, i.e., the communication stragglers (Wang et al., 2019; Wang and Joshi, 2021; Wang et al., 2020b). In particular, the celebrated “MATCHA” algorithm was proposed in (Wang et al., 2019) to mitigate the communication stragglers. Note that although MATCHA also has a link-control procedure, it differentiates from our proposed DSGD-AAU algorithm fundamentally. Specifically, DSGD-AAU adaptively selects the fastest computational workers for each iteration to mitigate the effect of slower computational workers, while MATCHA optimizes runtime by tuning the frequency of inter-node communication, potentially reducing the impact of slower communication workers, which leads DSGD-AAU to be an asynchronous algorithm and MACTHA to be a synchronous one. Furthermore, in contrast to MATCHA, our DSGD-AAU leverages the bounded-connectivity time to theoretically show the convergence rather than using the conventional spectral radius as in (Lian et al., 2017, 2018). These differences highlight that DSGD-AAU and MATCHA are designed to address distinct challenges in the field.
8. Conclusion
@@
Appendix A Additional Discussions on Pathsearch
@@
Algorithm 3 Pathsearch: logical view of arbitrary worker
Appendix B Proofs of Main Result
In this section, we provide proofs of the theoretical results presented in the paper, i.e., the convergence for DSGD-AAU after a large number of iterations (Theorem 4.1).
B.1. Proof of Theorem 4.1
@@ Appendix B Proofs of Main Result
and let ${\Phi}_{k,s}(i,j)$ be the $i$-th row $j$-th column element of $\boldsymbol{\Phi}_{k:s}$. Thus, with fixed $\eta$, ${\mathbf{w}}_{i}(k)$ can be expressed in terms of $\boldsymbol{\Phi}$ as
${\mathbf{w}}_{i}(k+1)$ $=\sum\limits_{j=1}^{N}\Phi_{k,1}(i,j){\mathbf{w}}_{j}(0)$
$\qquad-\eta\sum\limits_{r=1}^{k}\sum\limits_{j=1}^{N}\Phi_{k,r}(i,j)g_{i}({\mathbf{w}}_{j}(r-1)).$
@@
(8) $\bar{{\mathbf{w}}}(k)=\frac{1}{N}\sum\limits_{j=1}^{N}{\mathbf{w}}_{j}(0)-\eta\sum\limits_{r=1}^{k}\sum\limits_{j=1}^{N}\frac{1}{N}g({\mathbf{w}}_{j}(r-1)),$
@@
(9) $\bar{{\mathbf{w}}}(k+1)=\bar{{\mathbf{w}}}(k)-\frac{\eta}{N}\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k)).$
Due to the gradient Lipschitz-continuous in Assumption 4.3, we have the following inequality with respect to $\bar{{\mathbf{w}}}(k)$ (Bottou et al., 2018)
@@
$+\frac{L}{2}\mathbb{E}\left[\|\bar{{\mathbf{w}}}(k+1)\!-\!\bar{{\mathbf{w}}}(k)\|^{2}\right]\!+\!\mathbb{E}[F(\bar{{\mathbf{w}}}(k))]$
@@
(10) $+\underset{C_{2}}{\underbrace{\frac{L}{2}\mathbb{E}\left[\|\bar{{\mathbf{w}}}(k+1)\!-\!\bar{{\mathbf{w}}}(k)\|^{2}\right]}}+\!\mathbb{E}[F(\bar{{\mathbf{w}}}(k))]\!-\!\eta\|\nabla F(\bar{{\mathbf{w}}}(k))\|^{2}.$
As a result, we need to bound $C_{1}$ and $C_{2}$ in (B.1). The term $C_{1}$ can be bounded as follows
@@
$-\frac{\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))\right\|^{2}+\frac{\eta}{2}\left\|\nabla F(\bar{{\mathbf{w}}}(k))\right\|^{2}$
@@
(11) $-\frac{\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))\right\|^{2}+\frac{\eta}{2}\left\|\nabla F(\bar{{\mathbf{w}}}(k))\right\|^{2},$
where $(a_{1})$ follows from (9) and that $\frac{1}{N}\sum\limits_{j=1}^{N}\nabla F_{j}(\bar{{\mathbf{w}}}(k))=\nabla F(\bar{{\mathbf{w}}}(k))$; $(a_{2})$ comes from a standard mathematical manipulation; $(a_{3})$ is based on the equation $\langle{\mathbf{a}},{\mathbf{b}}\rangle=\frac{1}{2}\left[\|{\mathbf{a}}\|^{2}+\|{\mathbf{b}}\|^{2}-\|{\mathbf{a}}-{\mathbf{b}}\|^{2}\right]$ and $\mathbb{E}[g_{j}({\mathbf{w}}_{j}(k))]=\nabla F_{j}({\mathbf{w}}_{j}(k))$ from Assumption 4.4; and $(a_{4})$ is due to the well triangle-inequality $\left\|\sum\limits_{i=1}^{N}{\mathbf{a}}_{i}\right\|^{2}\leq N\sum\limits_{i}^{N}\left\|{\mathbf{a}}_{i}\right\|^{2}.$
@@
$D_{1}{\leq}$ $\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\Big{[}\|{\mathbf{w}}_{j}(k)-\bar{{\mathbf{w}}}(k)\|^{2}\Big{]},$
where the inequality comes from Assumption 4.3. Substituting $D_{1}$ into $C_{1}$ yields
$C_{1}$ $\leq\frac{\eta L^{2}}{2N}\sum\limits_{j=1}^{N}\mathbb{E}\left\|{\mathbf{w}}_{j}(k)\!-\!\bar{{\mathbf{w}}}(k)\right\|^{2}\!-\!\frac{\eta}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))\right\|^{2}$
(12) $\qquad\qquad+\frac{\eta}{2}\left\|\nabla F(\bar{{\mathbf{w}}}(k))\right\|^{2}.$
@@
=$\overset{(c_{1})}{=}\frac{\eta^{2}L}{2N^{2}}\Bigg{(}\mathbb{E}\left[\left\|\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))-\nabla F_{j}({\mathbf{w}}_{j}(k))\right\|^{2}\right]$
$\qquad+\left\|\mathbb{E}\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))\right\|^{2}\Bigg{)}$
(13) $\overset{(c_{2})}{\leq}\frac{\eta^{2}L}{2N}\sigma^{2}_{L}+\frac{\eta^{2}L}{2N^{2}}\mathbb{E}\left\|\sum\limits_{j=1}^{N}g_{j}({\mathbf{w}}_{j}(k))\right\|^{2},$
@@
Substituting $C_{1}$ and $C_{2}$ into the inequality (B.1), we have